### v3

# Efficient Analysis of Latent Spaces in Heterogeneous Networks

Abstract
This work proposes a unified framework for efficient estimation under latent space modeling of heterogeneous networks. We consider a class of latent space models that decompose latent vectors into shared and network-specific components across networks. We develop a novel procedure that first identifies the shared latent vectors and further refines estimates through efficient score equations to achieve statistical efficiency. Oracle error rates for estimating the shared and heterogeneous latent vectors are established simultaneously. The analysis framework offers remarkable flexibility, accommodating various types of edge weights under general distributions.
Keywords: Network, Latent space model, Data integration, Low rank, Heterogeneity.
1 Introduction
In recent years, analyses of multiview or multiplex networks (Kivelä et al.,, 2014) open up new opportunities for understanding complex systems across diverse scientific endeavors (Salter-Townshend and McCormick,, 2017), including social science (Lazega,, 2001; Banerjee et al.,, 2013), microbiome (Gould et al.,, 2018), and neuroscience (Wen et al.,, 2022). In these applications, it is common to observe multiple networks sharing a common set of nodes, with each edge representing specific types of relationships between node pairs. Studying multiple networks enables a unified understanding of multifaceted and interconnected systems.
Latent space modeling has emerged as an important approach in network analysis (Hoff et al.,, 2002; Matias and Robin,, 2014). In this framework, each node is mapped to a vector in a latent space, and the connectivity strength between two nodes is parametrized through a function between two corresponding latent vectors. While latent spaces can be discrete, they are more commonly referred to as stochastic block models and exhibit unique properties from discreteness (Holland et al.,, 1983). Throughout this paper, we focus on continuous latent vectors and consider latent spaces as Euclidean spaces following the convention in the literature (Hoff et al.,, 2002).
For multiple or multilayer networks, various developments under latent space modeling have been proposed, including Bayesian approaches (Hoff,, 2011; Salter-Townshend and McCormick,, 2017), and frequentist approaches (Arroyo et al.,, 2021; Jones and Rubin-Delanchy,, 2020; Nielsen and Witten,, 2018; Zheng and Tang,, 2022; Zhang et al.,, 2020; MacDonald et al.,, 2022). Our work aligns with the latter, focusing on estimating fixed latent vectors.
Existing studies have consistently shown that multiple networks often share underlying structures while simultaneously exhibiting significant network-specific heterogeneity (Wang et al.,, 2019; Zhang et al.,, 2020; Arroyo et al.,, 2021; Chen et al.,, 2022; He et al.,, 2025; MacDonald et al.,, 2022). When analyzing a collection of heterogeneous networks, leveraging their shared information is crucial for enhancing statistical efficiency, whereas accurately characterizing unique features is equally important. The coexistence of shared and unique patterns calls for analysis capable of disentangling the intertwined structural features and still fully exploiting statistical efficiency. Recently, MacDonald et al., (2022) introduced a general class of latent space models for multiple networks that decompose latent vectors into shared and network-specific components. Despite flexibility of the models, fundamental statistical limits under them remain unclear. Specifically, a critical statistical question for jointly analyzing multiple networks is: Can pooling heterogeneous networks improve the efficiency of estimating latent vectors, and if so, how can such gains be effectively achieved?
To address this, we develop a unified framework of efficient analysis under the latent space modeling of multiple heterogeneous networks. We propose a novel strategy that first hunts for the shared latent space across networks and then applies a refinement to achieve oracle efficiency. In this way, challenges from identifying shared space and achieving efficiency can be tackled separately, enhancing analytical flexibility. To theoretically quantify the efficiency gain, we develop novel techniques that disentangle the distinct oracle estimation error rates of shared and network-specific spaces through analyzing complex efficient score equations. Our proposed framework is versatile and encompasses a broad range of edge weight distributions.
The rest of the paper is organized as follows. Section 2 introduces the model and discusses statistical challenges of efficient estimation. Section 3 presents our proposed procedures, and Section 4 establishes the corresponding estimation error rates. Sections 5 and 6 present simulations and data analysis, respectively. Section 7 discusses several extensions of the proposed framework. Additional details and proofs are deferred to the Supplementary Material.
We summarize notations used below. Given two sequences of real numbers $\{g_{n}\}$ and $\{h_{n}\}$, $g_{n}=O(h_{n})$ and $g_{n}\lesssim h_{n}$ represent $|g_{n}|\leqslant c_{1}|h_{n}|$ for a constant $c_{1}$, $g_{n}\ll h_{n}$ and $g_{n}=o(h_{n})$ mean $\lim_{n\to\infty}g_{n}/h_{n}=0$, and $g_{n}\gg h_{n}$ means $\lim_{n\to\infty}h_{n}/g_{n}=0$. For a sequence of random variables $\{X_{n}\}$, write $X_{n}=O_{p}(g_{n})$ if for any $\epsilon>0$, there exists finite $M>0$ such that $\sup_{n}\Pr(|X_{n}/g_{n}|>M)<\epsilon$; write $X_{n}=o_{p}(g_{n})$ if for any $\epsilon>0$, $\lim_{n\to\infty}\Pr(|X_{n}/g_{n}|>\epsilon)=0$. For $x=(x_{i})_{i=1}^{n}$ and $y=(y_{i})_{i=1}^{n}\in\mathbb{R}^{n}$, $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}y_{i}$, $\|x\|_{2}=\sqrt{\langle x,x\rangle}$, and $\|x\|_{\infty}=\max_{1\leqslant i\leqslant n}|x_{i}|$. For $X=(x_{ij})_{1\leqslant i\leqslant n,1\leqslant j\leqslant m}\in\mathbb{R}^{n\times m}$, $\|X\|_{\mathrm{F}}=\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{m}x_{ij}^{2}}$, $\|X\|_{\operatorname{op}}=\sup_{\|v\|_{2}=1}\|Xv\|_{2}$, and $\|X\|_{2\to\infty}=\sup_{\|v\|_{2}=1}\|Xv\|_{\infty}$. Let $\mathrm{I}_{k}$ denote $k\times k$ identity matrix, and $\mathcal{O}(k)=\left\{Q\in\mathbb{R}^{k\times k}:QQ^{\top}=\mathrm{I}_{k}\right\}$ represent the set of $k\times k$ orthogonal matrices.
2 Model for heterogeneous networks
2.1 Latent space model with shared factors
This paper focuses on multiple networks where each network is observed on a common set of $n$ codes without inter-layer connections. Suppose we observe $T$ undirected networks on a common set of $n$ nodes, represented by $n\times n$ adjacency matrices $\{\mathbf{A}_{t}=(A_{t,ij})_{1\leqslant i,j\leqslant n}:t=1,\ldots,T\}$. For weighted networks, $A_{t,ij}$ is the weight value, which could be continuous or a count number. For unweighted networks, $A_{t,ij}$ is binary indicating that the connection between nodes $i$ and $j$ exists or not.
To model interactions between nodes, we adopt the popular latent space modeling framework (Hoff et al.,, 2002). Assume there exist latent vectors $y_{t,i}\in\mathbb{R}^{d_{t}}$ for $i=1,\ldots,n$ and $t=1,\ldots,T$ such that conditioning on $y_{t,i}$’s, network edges $A_{t,ij}$ are independently generated. To flexibly model different types of weight values, we allow $A_{t,ij}$ to be generated from a generic parametric distribution $p(\,\cdot\mid\theta\,)$ characterized by a parameter $\theta$, i.e.,
$A_{t,ij}=A_{t,ji}\ \sim\ p(\,\cdot\mid\Theta_{t,ij}\,),\quad\quad 1\leqslant i\leqslant j\leqslant n,\ 1\leqslant t\leqslant T,$ (1)
independently, where $\Theta_{t,ij}=\langle y_{t,i},y_{t,j}\rangle$ models the interaction effect between nodes $i$ and $j$ similarly to the popular low-rank models in the literature (Hoff et al.,, 2002; Ma et al.,, 2020). Examples of $p(\,\cdot\mid\theta\,)$ include Bernoulli, Gaussian, and Poisson distributions, commonly used to model edge weights that are binary, continuous, and counts, respectively. Model (1) posits a common $p(\,\cdot\mid\theta\,)$ across networks for notational simplicity. Our proposed analysis framework can be readily generalized when networks follow different types of distributions.
For each node $i$, we assume that the latent vectors $\{y_{t,i}:t=1,\ldots,T\}$ may include shared and heterogeneous components over $T$ networks following MacDonald et al., (2022). Specifically, $y_{t,i}=(z_{i}^{\top},w_{t,i}^{\top})^{\top}\in\mathbb{R}^{k+k_{t}}$, where $k+k_{t}=d_{t}$, $z_{i}\in\mathbb{R}^{k}$ represents a shared component over $T$ networks, and $w_{t,i}\in\mathbb{R}^{k_{t}}$ represents a heterogeneous component that may differ across $1\leqslant t\leqslant T$. Then $\Theta_{t,ij}=\langle y_{t,i},y_{t,j}\rangle=\langle z_{i},z_{j}\rangle+\langle w_{t,i},w_{t,j}\rangle.$
Let $\Theta_{t}=(\Theta_{t,ij})_{1\leqslant i,j\leqslant n}\in\mathbb{R}^{n\times n}$ denote the parameter matrix for the $t$-th network. We define $Y_{t}=[Z,\,W_{t}]\in\mathbb{R}^{n\times d_{t}},$ $Z=\left[z_{1},\ldots,z_{n}\right]^{\top}\in\mathbb{R}^{n\times k},$ and $W_{t}=\left[w_{t,1},\ldots,w_{t,n}\right]^{\top}\in\mathbb{R}^{n\times k_{t}}.$ Then our model has an equivalent low-rank matrix representation after link transformation
$\mu^{-1}\{\mathbb{E}(\mathbf{A}_{t})\}=\Theta_{t}=Y_{t}Y_{t}^{\top}=ZZ^{\top}+W_{t}W_{t}^{\top},\hskip 15.00002pt1\leqslant t\leqslant T,$ (2)
where $\mu(\cdot)$ denotes the unique invertible function linking $\theta$ and variable expectation under $p(\,\cdot\mid\theta\,)$, and it operates entrywisely on a matrix argument in (2).
Remark 1
The core model (1) is flexible and can be readily adapted to accommodate diverse network properties. For example, to allow flexible node degrees, an overall degree factor varying with respect to $(n,T)$ can be added. Also, (1) can be augmented with an extra probability component through mixture modeling, enabling features such as link missingness or disconnections in weighted networks. Our developments can be similarly established under these generalized models, exemplifying the broader applicability of our framework. Detailed results are provided in Sections J–L of the Supplementary Material.
2.2 Identifiability
By (2), the parameters $Z$ and $W_{t}$ can only be identified up to orthogonal transformation, i.e., $ZZ^{\top}=ZQQ^{\top}Z^{\top}$ and $W_{t}W_{t}^{\top}=W_{t}Q_{t}Q_{t}^{\top}W_{t}^{\top}$ for any $Q\in\mathcal{O}(k)$ and $Q_{t}\in\mathcal{O}(k_{t})$. We say that parameters $Z$ and $W_{t}$ in the model (2) are identifiable up to orthogonal group transformation if, for any two sets of parameters $\{Z,W_{1},\ldots,W_{T}\}$ and $\{Z^{\prime},W_{1}^{\prime},\ldots,W_{T}^{\prime}\}$ yielding the same model (2), there exist $Q\in\mathcal{O}(k)$ and $Q_{t}\in\mathcal{O}(k_{t})$ such that $Z=Z^{\prime}Q$ and $W_{t}=W^{\prime}_{t}Q_{t}$ for $1\leqslant t\leqslant T$. We next establish a sufficient condition for the identifiability of the model (2).
Proposition 1
The model (2) is identifiable up to orthogonal group transformation if
  • (i)
    for any $1\leqslant t\leqslant T$, the columns of $Y_{t}=[Z,W_{t}]$ are linearly independent;
  • (ii)
    there exist $1\leqslant t<s\leqslant T$ such that the columns of $Z,W_{t},W_{s}]$ are linearly independent.
Remark 2
Proposition 1 is adapted from MacDonald et al., (2022) but relaxes the connectivity assumption of the inter-layer graphs. Given (i) and (ii) in Proposition 1, the column span of $Z$ equals the intersection of all column spans of $Y_{t}$ across $1\leqslant t\leqslant T$.
2.3 Efficient analysis of latent spaces and challenges
Estimating latent spaces can reveal intrinsic connectivity patterns underlying noisy networks. Analyzing estimation efficiency provides crucial insights into the quality of obtained representations and is vital for downstream tasks. While MacDonald et al., (2022) introduced the general model, their theoretical analysis was constrained to Gaussian distribution with identity link. There still lacks comprehensive understanding into efficiency gain from pooling heterogeneous networks under general distributions.
For longitudinal networks with baseline node degrees varying over time, He et al., (2025) established semiparametric efficient estimation for shared latent spaces. Specifically, He et al., (2025) showed that the aggregated squared estimation error of shared latent factors can reach the oracle rate that is inverse proportional to the number of networks $T$, whereas the estimation errors of baseline nuisance parameters do not decrease with respect to $T$.
For estimating the shared factor $Z$ under (2), we expect a similar semiparametric oracle error rate. As an illustration, consider an oracle scenario when heterogeneous (nuisance) parameters $w_{t,i}$’s are known. In this case, each shared (target) parameter $z_{i}$ is measured by $nT$ independent edges $\{A_{t,ij}:j=1,\ldots,n;t=1,\ldots,T\}$. The oracle squared estimation error rate of each latent vector $z_{i}$ is expected to be of the order of $O_{p}(k/(nT))$, the ratio between the parameter number and the effective sample size (Portnoy,, 1988). Thus the aggregated oracle estimation error of $n$ latent vectors $Z$ is expected to be $O_{p}(n\times k/(nT))=O_{p}(1/T)$ under fixed $k$. Theoretically, achieving this oracle error rate requires approximating Cramér-Rao lower bound for high-dimensional $Z$; more discussions on the notion of efficiency are given in Remark D.4 of the Supplementary Material. Nevertheless, to establish the result, technical challenges similar to those in He et al., (2025) persist.
  1. 1.
    Oracle error rates for estimating target and nuisance parameters are different, but their errors are entangled in the analysis. Intuitively, each nuisance $w_{t,i}$ is measured by $n$ independent edges: $\left\{A_{t,ij}:j=1,\ldots,n\right\}$, so the oracle squared error rate of each $w_{t,i}$ is expected to be $O_{p}(1/n)$, larger than $O_{p}(1/(nT))$ of each $z_{i}$. In theoretical analysis, it is difficult to separate the error of $z_{i}$ from that of $w_{t,i}$ due to several reasons. One is that the individual components $w_{t,i}$’s encode both node-specific heterogeneity over $i$ and network-specific heterogeneity over $t$. As a result, classical partial likelihood approach (Andersen and Gill,, 1982) cannot directly remove the heterogeneous factors. Moreover, when the link function in (2) is non-linear, such as the logistic or exponential function, semiparametric oracle rates for $Z$ cannot be easily obtained by spectral-based analyses (Arroyo et al.,, 2021) or analysis under Gaussian distribution (MacDonald et al.,, 2022).
  2. 2.
    Target parameters are unidentifiable. As argued in Section 2.2, the shared (target) latent space $Z$ is unidentifiable in the Euclidean space. Actually, the intrinsic space of $Z$ is a quotient set with the equivalence relation up to the orthogonal group transformation. To evaluate the discrepancy between a target matrix $Z^{\star}$ and an estimate $\hat{Z}$, a natural distance is $\mathrm{dist}(\hat{Z},Z^{\star})=\min_{Q\in\mathcal{O}(k)}\|\hat{Z}-Z^{\star}Q\|_{\mathrm{F}}$, instead of the Euclidean distance.
  3. 3.
    Both target and nuisance parameters are high-dimensional. Specifically, under (2), the dimension of $Z$ increases with respect to the network size $n$, and the total number of parameters in $W_{1},\ldots,W_{T}]$ grows with respect to both $n$ and $T$.
On the other hand, our model (2) posits unique challenges and properties, making existing analyses inapplicable and necessitating new developments of methodology and theory. We summarize critical distinctions into four key aspects.
  1. 1.
    Cross-network identification. Identifying the shared latent space under (2) requires joint information across multiple networks, as shown in Proposition 1. This fundamentally distinguishes from He et al., (2025) that only needs a single network to identify the target latent space. Therefore, new procedures are required to effectively extract the shared latent space, and theoretical analysis must properly integrate multiple networks together.
  2. 2.
    Intricate factor interrelationships. The relationship between the shared (target) space $Z$ and individual (nuisance) spaces $W_{t}$’s exhibits unprecedented complexity. In He et al., (2025), the target and nuisance parameters are latent factors and baseline degrees, respectively, which characterize intrinsically orthogonal information that can enhance theoretical analysis. In contrast, factor interrelationships in (2) defy simple characterizations, which will be further elaborated in Remark 4. This complexity greatly amplifies the challenge of separating estimation errors across different $1+T$ factor groups.
  3. 3.
    Unidentifiable nuisance parameters. Similar to the target $Z$, the network-specific nuisance parameters $W_{t}$’s are unidentifiable in Euclidean space. This contrasts sharply with He et al., (2025), which worked with identifiable nuisance degree parameters. Consequently, we develop entirely novel proof techniques to tackle the unidentifiability of nuisance parameters and establish uniform error control across them.
  4. 4.
    Distributional generalization. The model in (1) accommodates a broad family of distributions, significantly extending beyond the Poisson edge weight distributions assumed in He et al., (2025). This added flexibility enhances the model’s applicability but also leads to additional analytical complexity.
3 Estimation of the latent spaces
In the following, we assume the observed data follow the model (2) with parameters $(Z,W_{t})=(Z^{\star},W_{t}^{\star})$ over $1\leqslant t\leqslant T$. We aim to estimate the underlying true parameters $Z^{\star}$ and $W^{\star}_{t}$ from the observed data. To address all the challenges discussed in Section 2.3, we propose a three-stage estimation procedure:
  • (a)
    for $1\leqslant t\leqslant T$, estimate concatenated latent factors $Y_{t}^{\star}=[Z^{\star},W_{t}^{\star}]$ up to an orthogonal group transformation in each $t$-th network individually;
  • (b)
    separate shared factors $Z^{\star}$ and heterogeneous factors ${W}_{t}^{\star}$ based on joint information across the $T$ estimates in Step (a) above;
  • (c)
    refine the estimators with likelihood information to achieve oracle efficiency.
As Step (a) analyzes each network individually, $Y_{t}^{\star}$ can only be identified up to an orthogonal group transformation $\mathcal{O}(d_{t})$ under (2). Consequently, column spaces of $Z^{\star}$ and $W_{t}^{\star}$ cannot be separated, necessitating a joint analysis across $T$ networks in Steps (b) and (c). As Section 2.2 suggests, $Z^{\star}$ and $W_{t}^{\star}$ can only be estimated up to factor-specific orthogonal group transformations $\mathcal{O}(k)$ and $\mathcal{O}(k_{t})$, respectively. For simplicity, we refer to estimating $Z^{\star}$ and $W_{t}^{\star}$ without explicitly emphasizing these group transformations when there is no ambiguity. In the following, we assume the latent dimensions $k$ and $k_{t}$’s are fixed and given, and discuss how to consistently estimate them in Section B of the Supplementary Material.
3.1 Individual estimation
We first estimate each $Y_{t}^{\star}$ up to an orthogonal group transformation, which can be achieved by examining each single network individually. For a single network, there have been extensive studies to estimate latent spaces for both random dot product models and latent space models with non-linear link function (Chatterjee,, 2015; Ma et al.,, 2020; Zhang et al.,, 2020). For the subsequent analysis in Sections 3.2 and 3.3, it suffices to obtain an estimator $\mathring{Y}_{t}$ satisfying $\mathrm{dist}^{2}(\mathring{Y}_{t},Y_{t}^{\star})$ being of the order of $O_{p}(1)$ up to logarithmic factors.
The desired error rate is consistent with existing results in the literature and may be achieved by various methods. In this work, we adopt a strategy that maximizes the likelihood function of $Y_{t}^{\star}$ via projected gradient descent for its simplicity and computational efficiency. The method generalizes that in Ma et al., (2020) for a binary network to models under a general edge distribution. As we consider a general link function in the model (2), different adjustments for links are needed are in both methodology and theory. We defer the details to Section A in the Supplementary Material. Notably, any other methods that can estimate $Y_{t}^{\star}$ with the desired error rate can also be used in our subsequent analysis.
3.2 Shared space hunting
Although estimating $Y_{t}^{\star}=[Z^{\star},W_{t}^{\star}]$ from each network is well-established, $Y_{t}^{\star}$ is only identifiable up to an unknown orthogonal transformation under a single-network model. As a result, $Z^{\star}$ cannot be directly obtained. To address this issue, we leverage joint information of $Y_{t}^{\star}$’s and develop a novel and efficient spectral method below. To motivate our proposed construction, we first introduce an algebraic result in an oracle scenario below.
Proposition 2
Consider an oracle scenario where
  • (i)
    we obtain estimators ${Y}_{t}\in\mathbb{R}^{n\times d_{t}}$ satisfying ${Y}_{t}{Y}_{t}^{\top}=Y_{t}^{\star}Y_{t}^{\star\top}$ for all $1\leqslant t\leqslant T$;
  • (ii)
    there exist $1\leqslant t<s\leqslant T$ such that columns of $Z^{\star},W_{t}^{\star},W_{s}^{\star}]$ are linearly independent.
Let $\mathcal{T}=\{(t,s):1\leqslant t<s\leqslant T\text{ and Assumption (ii) holds for }(t,s)\}$. For $(t,s)\in\mathcal{T}$,
$Z^{\star}Z^{\star\top}=Y_{t,-s}\,V_{t,s}\,V_{t,s}^{\top}\,Y_{t,-s}^{\top}/2,$ (3)
where we let $Y_{t,-s}=[Y_{t},\,-Y_{s}]\in\mathbb{R}^{n\times(d_{t}+d_{s})}$, and $V_{t,s}\in\mathbb{R}^{(d_{t}+d_{s})\times k}$ is a matrix whose columns can be any set of basis of the null space of $Y_{t,s}=[Y_{t},Y_{s}]\in\mathbb{R}^{n\times(d_{t}+d_{s})}$.
Assumption (i) in Proposition 2 posits an oracle situation where we can accurately estimate each $Y_{t}^{\star}$ individually up to the orthogonal group transformation $\mathcal{O}(d_{t})$. Assumption (ii) in Proposition 2 is the same as Assumption (ii) in Proposition 1. Eq. (3) provides the theoretical foundation for identifying $Z^{\star}$, which is unachievable when only investigating each $Y_{t}$ individually. However, (3) may not be directly applicable to estimating $Z^{\star}$ in practice because an estimator $\mathring{Y}_{t}$ obtained in Step (a) may not exactly satisfy $\mathring{Y}_{t}\mathring{Y}_{t}^{\top}=Y_{t}^{\star}Y_{t}^{\star\top}$, and $\mathcal{T}$ relies on unknown true parameters.
To address these issues, we develop practical procedures based on spectral information of the estimates $\mathring{Y}_{t}$’s in Step (a). We first estimate $\mathcal{T}$. Let $\sigma_{i}(B)$ denote the $i$-th largest singular value of a matrix $B$, and define $\mathcal{R}_{i,j}(B)=\sigma_{i}(B)/\sigma_{j}(B)$. Then we have
$\sigma_{k+k_{t}+k_{s}}\big{(}\,[Y_{t}^{\star},Y_{s}^{\star}]\,\big{)}\ \begin{cases}\ =0,&\ (t,s)\not\in\mathcal{T};\\ \ \neq 0,&\ (t,s)\in\mathcal{T}.\end{cases}$
When the estimates $\mathring{Y}_{t}$’s are close to true $Y_{t}^{\star}$’s, we expect that for $(t,s)\not\in\mathcal{T}$, $\sigma_{k+k_{t}+k_{s}}(\mathring{Y}_{t,s})$ with $\mathring{Y}_{t,s}=[\mathring{Y}_{t},\mathring{Y}_{s}]$ is close to 0, and then $\mathcal{R}_{1,k+k_{t}+k_{s}}(\mathring{Y}_{t,s})$ would diverge to infinity quickly. Based on this idea, we construct an estimated set $\hat{\mathcal{T}}$ by letting $(t,s)\in\hat{\mathcal{T}}$ if the corresponding ratio $\mathcal{R}_{1,k+k_{t}+k_{s}}(\mathring{Y}_{t,s})$ is below an appropriate threshold.
Moreover, as (3) cannot hold exactly due to the estimation error in $\mathring{Y}_{t}$, we stabilize the result by investigating the averaged terms over index pairs in $\hat{\mathcal{T}}$, i.e., we define
$\mathring{F}=\frac{1}{2|\hat{\mathcal{T}}|}\sum_{(t,s)\in\hat{\mathcal{T}}}\mathring{Y}_{t,-s}\,\mathring{V}_{t,s}\,\mathring{V}_{t,s}^{\top}\,\mathring{Y}_{t,-s}^{\top},$ (4)
where $\mathring{V}_{t,s}\in\mathbb{R}^{(d_{t}+d_{s})\times k}$ is a matrix consisting of the right null singular vectors of $\mathring{Y}_{t,s}$, and $\mathring{Y}_{t,-s}=[\mathring{Y}_{t},-\mathring{Y}_{s}].$ By (3), we estimate $Z^{\star}$ by squared root of the top $k$ eigen-decomposition of $\mathring{F}$, represented as $\mathcal{S}_{k}(\mathring{F})$, which is formally defined in (A.1) of the Supplementary Material.
3.3 Likelihood-based refinement
The spectral method in Section 3.2 is computationally efficient and could be used under weak assumptions on data distribution. However, when the data distribution is available, the spectral method may not be optimal for estimating the shared space $Z^{\star}$ in the sense of statistical efficiency discussed in Section 2.3, as the joint likelihood information is not fully exploited. We next propose a likelihood-based refinement procedure and theoretically demonstrate its optimality in Section 4.
Under the model (1), the log-likelihood function with respect to $Z$ and $W_{t}$’s satisfies
$\ell(Z,W)=\sum_{t=1}^{T}\sum_{1\leqslant i\leqslant j\leqslant n}l\big{(}\langle z_{i},z_{j}\rangle+\langle w_{t,i},w_{t,j}\rangle;A_{t,ij}\big{)},$ (5)
where $l(\theta;x)=\log p(\,x\mid\theta\,)$ and $W=[W_{1},\ldots,W_{T}]$. Our refinement procedure first conducts the projected gradient descent (Chen and Wainwright,, 2015), primarily using first-order derivatives of the likelihood function that can be efficiently computed. Specifically, the algorithm descends parameters along their gradient directions with pre-specified step sizes and then projects updated estimates to pre-specified constraint sets for parameters. The detailed steps are summarized in Algorithm 2 with notations explained in Section 3.4. Let $\check{Z}$ and $\check{W}_{t}$’s denote the estimates after sufficient iterations of the projected gradient descent.
Then the refinement procedure further constructs a second-order update, which utilizes efficient influence function and provides convenience for theoretically establishing desired statistical error rate. More discussions on its theoretical usefulness are provided in Remark D.6 of the Supplementary Material. For the ease of introducing formula below, we denote a vectorization of all the entrywise parameters of $Z,W_{1},\ldots,W_{T}]$ as
$v=\left[z_{1}^{\top},\ldots,z_{n}^{\top},w_{1,1}^{\top},\ldots,w_{1,n}^{\top},\ldots,w_{T,1}^{\top},\ldots,w_{T,n}^{\top}\right]^{\top}\in\mathbb{R}^{n(k+k_{\operatorname{sum}})\times 1},$ (6)
where $k_{\operatorname{sum}}=k_{1}+\cdots+k_{T}$, and similarly define $\check{v}$ as the vectorization of $\check{Z},\check{W}_{1},\ldots,\check{W}_{T}]$. Then $\ell(Z,W)=\ell(v)$, and let $\dot{\ell}(v)$ denote the partial derivative of $\ell(v)$ with respect to $v$. We construct a second-order update as
$\hat{v}=\check{v}+I(\check{v})^{+}\dot{\ell}(\check{v}),$ (7)
where $B^{+}$ represents the pseudo-Moore-Penrose inverse of a matrix $B$, which is uniquely defined, and $I(v)=\mathbb{E}\{\dot{\ell}(v)\dot{\ell}(v)^{\top}\}$ is the Fisher information matrix of $v$ with expectation taken under the model (1) with parameters $v$. An analytical formula for $I(v)$ is derived for practical use in Section D.5.3 of the Supplementary Material. The final estimators $\hat{Z}$ and $\hat{W}_{t}$’s are obtained through realigning $\hat{v}$ into matrices.
Remark 3
Our proof shows $I(\check{v})$ in (7) is singular, necessitating the use of pseudo inverse in (7). Intuitively, this singularity arises from the unidentifiability of latent factors $Z$ and $W_{t}$ (Little et al.,, 2010). It makes the theoretical analysis challenging and conclusions for classical one-step estimator (Van der Vaart,, 2000) or a single network Xie and Xu, (2023) inapplicable; more details are discussed in Remark D.5 of the Supplementary Material. Moreover, (7) is simultaneously constructed for $1+T$ groups of latent factors $Z,W_{1},\ldots,W_{T}]$. This markedly differs from He et al., (2025) that constructs one-step estimator for shared latent factors only and requires network-specific parameters to be identifiable. As a result, our theoretical analysis pioneers a new approach that jointly analyzes all the parameters in (6) and tackles inherent singularity of the information matrix from multiple groups of latent factors. More details are given in Section D.5.4 of the Supplementary Material.
3.4 Algorithm details
Algorithms 1 and 2 summarize the construction of the space-hunting estimator in Section 3.2 and the likelihood-refinement estimator in Section 3.3, respectively. We next introduce detailed notations and discuss the choice of hyperparameters in the algorithms.
Input: Individual estimates: $\mathring{Y}_{t}$ for $1\leqslant t\leqslant T$. Parameter: $\tau_{1}$ (threshold).
Output: $\mathring{Z}$ and $\mathring{W}_{t}$ for $1\leqslant t\leqslant T$.
1
2Let $\hat{\mathcal{T}}$ be an empty set.
3for $1\leqslant t<s\leqslant T$ do
4   Let $\mathring{Y}_{t,s}=[\mathring{Y}_{t},\,\mathring{Y}_{s}]$.
5 if $\mathcal{R}_{1,\,k+k_{t}+k_{s}}(\mathring{Y}_{t,s})\leqslant\tau_{1}$ then
[…]
7     Construct $\mathring{V}_{t,s}$ as the matrix whose columns consist of the right singular vectors of $\mathring{Y}_{t,s}$ corresponding to its $k$ smallest singular values.
8   end if
9
10 end for
11
12Let $\mathring{Z}=\mathcal{S}_{k}(\mathring{F})$ for $\mathring{F}$ in (4) and $\mathcal{S}_{k}(\cdot)$ defined in (A.1).
13Let $\mathring{W}_{t}=\mathcal{S}_{k_{t}}(\mathring{Y}_{t}\mathring{Y}_{t}^{\top}-\mathring{F})$ for $1\leqslant t\leqslant T$, where $\mathcal{S}_{k_{t}}(\cdot)$ is defined similarly to (A.1).
Algorithm 1 Spectral-based shared space hunting.
Input: Data: $\mathbf{A}_{1},\ldots,\mathbf{A}_{T}\in\mathbb{R}^{n\times n}$. Initial estimates: $\mathring{Z}$ and $\mathring{W}_{t}$ for $1\leqslant t\leqslant T$.
     Parameters: $\eta_{Z}$, $\eta_{W}$ (step sizes), $R$ (number of iterations),
          $\mathcal{C}_{Z},\mathcal{C}_{W_{t}}$ for $1\leqslant t\leqslant T$ (constraint sets for projection).
1
Output: $\hat{Z}$ and $\hat{W}_{t}$ for $1\leqslant t\leqslant T$.
2
3Let $Z^{0}=\mathring{Z}$ and $W_{t}^{0}=\mathring{W}_{t}$ for $1\leqslant t\leqslant T$.
4for $r=0,\ldots,R-1$ do
5
6  Let $\tilde{Z}^{r+1}=Z^{r}+{\eta_{Z}}\,\dot{\ell}_{Z}(Z^{r},W^{r})$ and $Z^{r+1}={\mathcal{P}_{\mathcal{C}_{Z}}}(\tilde{Z}^{r+1})$.
7  Let $\tilde{W}_{t}^{r+1}=W_{t}^{r}+\eta_{W}\,\dot{\ell}_{W_{t}}(Z^{r},W^{r})$ and $W_{t}^{r+1}=\mathcal{P}_{\mathcal{C}_{W_{t}}}(\tilde{W}_{t}^{r+1})$ for $1\leqslant t\leqslant T$.
8 end for
9
10Construct $\hat{v}$ by (7) with $(\check{Z},\check{W})=(Z^{R},W^{R})$, and let $\hat{Z},\hat{W}$] be its matrix version.
Algorithm 2 Likelihood-based refinement.
In Algorithm 1: The threshold $\tau_{1}$ is chosen to estimate $\mathcal{T}$. As Section 3.2 suggests, the choice of $\tau_{1}$ depends on statistical properties of the input initial estimates $\mathring{Y}_{t}$. In this paper, we construct $\mathring{Y}_{t}$ by the method in Section 3.1, whose estimation error is established in Section A.2 of the Supplementary Material. To achieve theoretical guarantee, it suffices to set the threshold to satisfy $1\ll\tau_{1}\lesssim\sqrt{\log n}$. Our numerical implementation sets $\tau_{1}=\sqrt{2\log n}$.
In Algorithm 2: During the gradient descent, we let $\dot{\ell}_{Z}(Z,W)\in\mathbb{R}^{n\times k}$ and $\dot{\ell}_{W_{t}}(Z,W)\in\mathbb{R}^{n\times k_{t}}$ represent the partial derivatives of $\ell(Z,W)$ with respect to matrices $Z$ and $W_{t}$, respectively. Let $\mathcal{P}_{\mathcal{C}}(\cdot)$ denote a projection operator given a pre-specified constraint set $\mathcal{C}$, and we consider two constraint sets
$\mathcal{C}_{Z}=\big{\{}Z\in\mathbb{R}^{n\times k}:\|Z\|_{2\to\infty}\leqslant M_{1}\big{\}}\quad\text{and}\quad\mathcal{C}_{W_{t}}=\big{\{}W_{t}\in\mathbb{R}^{n\times k_{t}}:\|W_{t}\|_{2\to\infty}\leqslant M_{1}\big{\}},$ (8)
for $Z$ and $W_{t}$, respectively. Eq. (8) constraints the two-to-infinity norms of $Z$ and $W_{t}$ to be bounded by $M_{1}$, which also corresponds to the constraints on the true parameters $(Z^{\star},W_{t}^{\star})$ in Condition 2 below. The projection steps and second-order update in Algorithm 2 are required for the convenience of the proof, and skipping them can yield equally good numerical performance. More details are discussed in Section G.2 of the Supplementary Material. In our numerical implementation, we use Barzilai-Borwein step sizes (Barzilai and Borwein,, 1988) and set $R=1000$. To establish theoretical results in Section 4, Condition 1 below summarizes the requirements on the hyperparameters in Algorithms 1 and 2.
Condition 1
Assume the hyperparameters in Algorithms 1 and 2 satisfy: (i) $1\ll\tau_{1}\lesssim\sqrt{\log n}$. (ii) Step sizes $\eta_{Z}=\eta/(nT)$ and $\eta_{W}=\eta/n$ for a sufficiently small constant $\eta>0$. (iii) Constraint sets $\mathcal{C}_{Z}$ and $\mathcal{C}_{W_{t}}$ are chosen as in (8). (iv) Number of iterations $R\gg\log(nT)$.
3.5 Related works
The proposed estimation scheme in Section 3 shares connections with, yet notably differs from, existing methods. The shared space hunting Algorithm 1 is developed based on spectral properties of latent factors. While spectral-based analysis is common in network studies (Rohe et al.,, 2011; Chatterjee,, 2015; Arroyo et al.,, 2021), our procedure utilizes unique properties of multiple networks to separate the shared and distinct latent spaces, and thus cannot be directly implied by existing methods. The likelihood refinement Algorithm 2 utilizes the projected gradient descent, which is known to be computationally efficient for maximizing the non-convex likelihood functions for both single network (Ma et al.,, 2020) and multiple networks (Zhang et al.,, 2020). Nevertheless, the oracle estimation error rates under our model cannot be directly attained through existing technical developments of those methods. To address the theoretical challenges, Algorithm 2 further constructs a novel second-order update estimator (7) for multiple heterogeneous networks. The construction is motivated from the classical one-step estimator, typically obtained through a linear approximation of score equation (Van der Vaart,, 2000). Similar idea has also been recently used to construct efficient estimators for latent factors in network analysis, including the random dot product model for a single network (Xie and Xu,, 2023) and a semiparametric longitudinal network model (He et al.,, 2025). However, our model (1) substantially differs from those models in that it embraces a general edge distribution and also allows distinct latent factors over multiple networks. The extended flexibility posits unique challenges in identifying shared spaces and obtaining efficient estimation errors, necessitating novel developments of efficient score updating formula and theory.
The model (1) is motivated from MacDonald et al., (2022), which similarly considers (1) and assumes that part of the latent vectors are shared across the networks, i.e., $y_{t,i}=(z_{i}^{\top},w_{t,i}^{\top})^{\top}$. The difference is that in their work $\Theta_{t,ij}=\kappa(y_{t,i},y_{t,j})$ with $\kappa(x,y)=x^{\top}\mathrm{I}_{p,q}y$, where $\mathrm{I}_{p,q}$ represents the block-diagonal matrix with $\mathrm{I}_{p}$ and $\mathrm{I}_{q}$ on the upper left and lower right blocks, respectively. Although a more general similarity function is proposed, their theoretical analysis is restricted to Gaussian distribution with identity link, which can be inappropriate for network data with binary edges or count edges. Alternatively, we will next establish theoretical results that encompass various exponential family distributions. Please also see Remark 4 for more discussions on the comparison of technical conditions and results.
Table 1: Comparison of different models and results: link functions and the corresponding estimation error rates of shared and individual parameters (ER-Shared and ER-Individual).
Model Link $\mu(\theta)$ ER-Shared ER-Individual Constraints on $(n,T)$
Arroyo et al., (2021) $\theta$ $O_{p}\biggr{(}\frac{1}{n}+\frac{1}{T}\biggr{)}$ - -
MacDonald et al., (2022) $\theta$ $O_{p}\biggr{(}\frac{1}{T}\biggr{)}$ $O_{p}(1)$ $T=o(n^{{1}/{2}})$
Zhang et al., (2020) $\operatorname{expit}(\theta)$ $O_{p}\biggr{(}1+\frac{1}{T}\biggr{)}$ - $T=O(n)$
He et al., (2025) $\exp(\theta)$ $O_{p}\biggr{(}\frac{1}{n}+\frac{1}{T}\biggr{)}$ $O_{p}(1)$ $\mathrm{polylog}(T)=O(n)$
This work General $O_{p}\biggr{(}\frac{1}{n}+\frac{1}{T}\biggr{)}$ $O_{p}(1)$ $\mathrm{polylog}(T)=O(n)$
(Error rates above are simplified to emphasize their orders with respect to $n$ and $T$ up to logarithmic factors and also rescaled to be comparable. “-” represents inapplicable information. See more details in Section F of the Supplementary Material.)
Besides the studied model (1), there are various other models extracting node-wise latent embeddings from multiple networks. One research line extends the random dot product graphs (Young and Scheinerman,, 2007; Athreya et al.,, 2018) to multilayer networks (Arroyo et al.,, 2021; Zheng and Tang,, 2022). These models typically corresponds to an identity link function in (2), and thus developments cannot directly address the challenges under models with general non-linear link functions. Another class of models adopt non-linear canonical link functions (Hoff et al.,, 2002; Ma et al.,, 2020) and develop extensions for multiple networks (Zhang et al.,, 2020; MacDonald et al.,, 2022; He et al.,, 2025). Our developments under (1) advance the literature by establishing a unified theoretical framework that can encompass a variety of general link functions and allow heterogeneous latent factors with complex relationships. We summarize properties and results under the related models in Table 1, where our theory will be formally established in Section 4. Notably, oracle error rate of the shared parameters in this paper is consistent with earlier work in Arroyo et al., (2021), MacDonald et al., (2022), and He et al., (2025).
Targeting at distinct properties and challenges, researchers have also proposed other models for multiple networks. Examples include multilayer stochastic block models for community detection (Han et al.,, 2015; Paul and Chen,, 2016; Lei and Lin,, 2023), and tensor-decomposition-based models for effective dimension reduction (Lyu et al.,, 2023; Zhang and Wang,, 2023), etc. Although these models may also incorporate shared and heterogeneous structures across multiple networks, their analytical objectives and challenges differ from our focus on estimating continuous node-wise latent embeddings and thus are not directly compared in Table 1; see more discussions in Section F of the Supplementary Material.
4 Theory for estimating latent spaces
In this section, we establish estimation error rates of Algorithms 1 and 2 in Theorems 1 and 2, respectively. For technical developments, we impose regularity assumptions on the true parameters in Condition 2 and edge-wise distribution in Condition 3 below.
Condition 2
For a matrix $X\in\mathbb{R}^{n\times m}$, let $\mathcal{G}(X)=X^{\top}X/n\in\mathbb{R}^{m\times m}$ denote the Gram matrix of $X$ standardized by its row number, and let $\sigma_{\operatorname{min}}(X)$ denote the minimum singular value of $X$. Assume that there exist positive constants $M_{1}$, $M_{2}$, and $M_{3}$ such that
  • (i)
    $\|Z^{\star}\|_{2\to\infty}\leqslant M_{1}$ and for any $1\leqslant t\leqslant T$, $\|W_{t}^{\star}\|_{2\to\infty}\leqslant M_{1}$;
  • (ii)
    for any $1\leqslant t\leqslant T$, $\sigma_{\operatorname{min}}(G_{t}^{\star})\geqslant M_{2}$ for $G_{t}^{\star}=\mathcal{G}([Z^{\star},W_{t}^{\star}])$;
  • (iii)
    there exist $1\leqslant t<s\leqslant T$ such that $\sigma_{\operatorname{min}}(G_{t,s}^{\star})\geqslant M_{3}$ for $G_{t,s}^{\star}=\mathcal{G}([Z^{\star},W_{t}^{\star},W_{s}^{\star}])$.
We describe Condition 2 on one set of true parameters $(Z^{\star},W_{t}^{\star})$ for convenience and clarity. It is equivalent to impose Condition 2 on the equivalence classes of $Z^{\star}$ and $W_{t}^{\star}$ that are identifiable up to orthogonal transformations. In particular, Condition 2 is satisfied for $Z^{\star}$ and $W_{t}^{\star}$ if and only if it is satisfied for $Z^{\star}Q$ and $W_{t}^{\star}Q_{t}$ with $Q\in\mathcal{O}(k)$ and $Q_{t}\in\mathcal{O}(k_{t})$ over $1\leqslant t\leqslant T$, since the matrix norms remain the same.
Condition 2 (i) implies that $\ell_{2}$-norms of the latent vector $z_{i}^{\star}$ and $w_{t,i}^{\star}$ are bounded over all $i$ uniformly. It constrains all the parameters to be in a bounded space, which is a prevalent mild assumption in high-dimensional network analysis (Ma et al.,, 2020; Zheng and Tang,, 2022). Condition 2 (ii) and (iii) impose constraints on the scaled Gram matrices of the latent factors. They imply that the columns of $Z^{\star},W_{t}^{\star}]$ (for any $t$) and $Z^{\star},W_{t}^{\star},W_{s}^{\star}]$ (for one pair of $\{t,s\}$) are linearly independent, ensuring that the true model parameters are identifiable by Proposition 1. For Condition 2 (ii), note that
$\mathcal{G}([Z^{\star},W_{t}^{\star}])=\frac{1}{n}\begin{pmatrix}Z^{\star\top}Z^{\star}&Z^{\star\top}W_{t}^{\star}\\ W_{t}^{\star\top}Z^{\star}&W_{t}^{\star\top}W_{t}^{\star}\end{pmatrix}$
is a two-by-two block matrix. Condition 2 (ii) can be satisfied by directly imposing conditions on each block separately, and similarly for Condition 2 (iii). In particular, we establish Proposition 3 below providing sufficient conditions on the sub-matrices for Condition 2.
Proposition 3
If there exists a constant $M_{4}>0$ such that,
  • (a)
    for any $1\leqslant t\leqslant T$, $\sigma_{\operatorname{min}}(Z^{\star\top}Z^{\star}/n)\geqslant M_{4}$ and $\sigma_{\operatorname{min}}(W_{t}^{\star\top}W_{t}^{\star}/n)\geqslant M_{4}$,
  • (b)
    for any $1\leqslant t\leqslant T$, $\|Z^{\star\top}W_{t}^{\star}/n\|_{\operatorname{op}}\leqslant M_{4}/4$,
  • (c)
    and there exist $1\leqslant t<s\leqslant T$ such that $\|W_{t}^{\star\top}W_{s}^{\star}/n\|_{\operatorname{op}}\leqslant M_{4}/4$,
then Condition 2 (ii) and (iii) are satisfied with $M_{2}$ and $M_{3}$ that are fixed functions of $M_{4}$.
Remark 4
MacDonald et al., (2022) studies a similar model with generalized inner product (Rubin-Delanchy et al.,, 2022). Their theoretical analysis focuses on edge-wise normal distribution with $\mu(\theta)=\theta$, and their imposed conditions imply that as $n\to\infty$,
$\|Z^{\star\top}W_{t}^{\star}/n\|_{\operatorname{op}}\to 0,\quad\quad\|W_{t}^{\star\top}W_{s}^{\star}/n\|_{\operatorname{op}}\to 0.$ (9)
Proposition 3 suggests that Condition 2 (ii) and (iii) can relax (9). As $\|Z^{\star\top}W_{t}^{\star}/n\|_{\operatorname{op}}=0$ means that the columns of $Z^{\star}$ and $W_{t}^{\star}$ are orthogonal, we refer to the decaying rate in (9) as “nearly orthogonal” conditions. Such restrictions on the model parameters could limit the expressive power of the models for practical use. Methodologically, their proposed penalty performs convex relaxation of the rank constraints for $Z$ and all $W_{t}$’s individually. This strategy, intuitively, may not effectively take the relationship between latent vectors into account. As a result, method in MacDonald et al., (2022) could lead to different interpretations of estimated latent vectors. See more numerical illustrations in Sections 5 and 6.
Condition 3
Let $\mathcal{X}=\{x\in\mathbb{R}:p(\,x\mid\theta\,)>0\}$ denote the support of $p(\,\cdot\mid\theta\,)$. Assume $l(\theta;x)$ in (5) belongs to the natural exponential family and satisfies the following conditions.
  • (i)
    For any fixed $x\in\mathcal{X}$, $l(\theta;x)$ is three times differentiable with respect to $\theta$, with its first to third derivatives with respect to $\theta$ denoted by $l^{\prime}(\theta;x)$, $l^{\prime\prime}(\theta;x)$, and $l^{\prime\prime\prime}(\theta;x)$, respectively. Moreover, there exist positive constants $\kappa_{1}$, $\kappa_{2}$, and $\kappa_{3}$ such that $\kappa_{1}\leqslant-l^{\prime\prime}(\theta;x)\leqslant\kappa_{2}$ and $|l^{\prime\prime\prime}(\theta;x)|\leqslant\kappa_{3}$ for any $x\in\mathcal{X}$ and $|\theta|\leqslant 2M_{1}^{2}$, where $M_{1}$ is given in Condition 2.
  • (ii)
    There exists a constant $L>0$ such that $\mathbb{E}|l^{\prime}(\Theta_{t,ij}^{\star};A_{t,ij})|^{m}\leqslant\mathrm{var}\{l^{\prime}(\Theta_{t,ij}^{\star};A_{t,ij})\}L^{m-2}m!/2$ for any $1\leqslant i\leqslant j\leqslant n$, $1\leqslant t\leqslant T$, and any integer $m\geqslant 2$, where $m!=m(m-1)\cdots 1$ represents the factorial of $m$.
Condition 3 can be naturally satisfied by most distributions under natural exponential families. For instance, $l(\theta;x)\propto\theta x-\log(1+\exp(\theta))$ for Bernoulli distribution, and $l(\theta;x)\propto\theta x-\exp(\theta)$ for Poisson distribution (Efron,, 2022). The assumption of natural exponential family is adopted here primarily for the ease of presentation and understanding. We show that generalizations under more general distributions can be established following similar arguments in Section I of the Supplementary Material.
Given the conditions, we first establish the estimation error rates for the shared space hunting algorithm in Section 3.2.
Theorem 1
Assume Conditions 13. Let $(\mathring{Z},\mathring{W})$ be the estimators obtained through Algorithm 1 with $\mathring{Y}_{t}$ in Theorem A.2 as initialization. For any constant $\varepsilon>0$, there exist positive constants $c_{\varepsilon}$ and $C_{\varepsilon}$ such that when $\log^{d_{\operatorname{max}}+2}(nT)/n\leqslant c_{\varepsilon}$ with $d_{\operatorname{max}}=\max_{1\leqslant t\leqslant T}d_{t}$,
$\Pr\left[\left\{\operatorname{dist}^{2}(\mathring{Z},Z^{\star})+\max_{1\leqslant t\leqslant T}\operatorname{dist}^{2}(\mathring{W}_{t},W_{t}^{\star})\right\}>C_{\varepsilon}\log^{2}(nT)\right]=O((nT)^{-\varepsilon}).$
Theorem 1 suggests that with high probability, the overall estimation error rate of $\operatorname{dist}^{2}(\mathring{Z},Z^{\star})$ and $\operatorname{dist}^{2}(\mathring{W}_{t},W_{t}^{\star})$ are of the order of $O(1)$ up to logarithmic factors. This aligns with the expected oracle rate for $\Theta_{t}^{\star}=Z^{\star}Z^{\star\top}+W_{t}^{\star}W_{t}^{\star\top}$ as discussed in Section 2.3.
Remark 5
The shared space hunting Algorithm 1 requires initial individual estimates $\mathring{Y}_{t}$ for ${Y}_{t}^{\star}$ over $1\leqslant t\leqslant T$. As discussed in Section 3.1, our analysis focuses on $\mathring{Y}_{t}$ obtained by Algorithm A.2 for the ease of discussion. The estimation error rates of the proposed $\mathring{Y}_{t}$’s are established in Theorem A.2. Nevertheless, Theorem 1 can be similarly established using other initial estimates $\mathring{Y}_{t}$, as long as they can achieve a similar rate to that in Theorem A.2.
Theorem 1 only provides an upper bound on the estimation error rates for shared and heterogeneous parameters simultaneously. Establishing the desired efficiency for each group of parameters separately is challenging, as discussed in Section 2.3. To facilitate the theoretical analysis, we make an additional assumption on the heterogeneous parameters.
Condition 4
Assume $\|\mathcal{G}(W^{\star})/T\|_{\operatorname{op}}\leqslant M_{5}$, where $M_{5}=M_{2}^{2}/(8M_{1}^{2})$ and the constants $M_{1}$ and $M_{2}$ are given in Condition 2.
For $W^{\star}=[W_{1}^{\star},\ldots,W_{T}^{\star}]$, $\mathcal{G}(W^{\star})$ consists of submatrices $W_{t}^{\star\top}W_{s}^{\star}/n$ for $1\leqslant t\leqslant s\leqslant T$. Similarly to Condition 2, Proposition 4 gives a sufficient condition for interpreting Condition 4 based on the submatrices of $\mathcal{G}(W^{\star})$.
Proposition 4
Condition 4 is satisfied if with $M_{5}=M_{2}^{2}/(8M_{1}^{2})$,
  • (a)
    for any $1\leqslant t\leqslant T$, $\|W_{t}^{\star\top}W_{t}^{\star}/n\|_{\operatorname{op}}\leqslant M_{5}T/2$;
  • (b)
    for any $1\leqslant t<s\leqslant T$, $\|W_{t}^{\star\top}W_{s}^{\star}/n\|_{\operatorname{op}}\leqslant M_{5}/2$.
We compare Proposition 4 with Proposition 3 to gain insights into Condition 4. First, (a) in Proposition 4 specifies an upper bound on the singular values of $W_{t}^{\star\top}W_{t}^{\star}/n$, which is compatible with the lower bound (a) in Proposition 3 when $T$ is sufficiently large. Second, (b) in Proposition 4 specifies an upper bound on the singular values of $W_{t}^{\star\top}W_{s}^{\star}/n$, which differs from (b) in Proposition 3, as $M_{5}$ depends on constants $(M_{1},M_{2})$ in Condition 2. Condition 4 is primarily a technical condition needed to facilitate the control of certain high-order error terms in the analysis. Intuitively, it ensures that the column spaces of $W_{t}^{\star}$ and $W_{s}^{\star}$ for most pairs $t\neq s$ are sufficiently different, making it easier to disentangle their respective estimation errors. We point out that the constraint on $M_{5}$ may not limit the practical use of our method. Numerical studies in Section 5 demonstrate its robust performance even in the extreme case where $W_{t}^{\star}$’s are identical. This empirical evidence suggests that the constraint on $M_{5}$ could potentially be relaxed under certain scenarios, which, however, is not pursued here to preserve the clarity and coherence of our presentation.
Remark 6
[…]
We next show that the likelihood-based refinement procedure can further improve the estimation error rate in Theorem 1.
Theorem 2
[…]
$\Pr\left[\operatorname{dist}^{2}(\hat{Z},Z^{\star})>C_{\varepsilon}\max\left\{\frac{1}{T},\ \frac{1}{n}\right\}\log^{8}(nT)\right]=$ $~O\big{(}(nT)^{-\varepsilon}\big{)},$ (10)
$\Pr\left[\max_{1\leqslant t\leqslant T}\operatorname{dist}^{2}(\hat{W}_{t},W_{t}^{\star})>C_{\varepsilon}\log(nT)\right]=$ $~O\big{(}(nT)^{-\varepsilon}\big{)}.$ (11)
Theorem 2 suggests that with a high probability, up to logarithmic factors, the estimation error for $W_{t}^{\star}$ is of the order of $O(1)$, whereas the estimation error for $Z^{\star}$ is of the order of
$O\left[\max\left\{\frac{1}{T},\,\frac{1}{n}\right\}\right].$ (12)
[…]
Remark 7
As Table 1 suggests, Theorem 2 not only aligns with the best rates reported in the compared works but also advances the literature by accommodating the most flexible range of link functions. While He et al., (2025) reported similar error rates for both shared and individual parameters, our work distinguishes itself by addressing unique and previously insurmountable challenges highlighted in Section 2.3. We develop a comprehensive set of analytical innovations including effective shared space identification validated by Theorem 1, and a novel joint one-step construction framework addressing the multiplex singularity issue of the efficient information matrix in Remark 3. These breakthroughs collectively establish a solid foundation for comprehensive analysis of heterogeneous networks, deepening our understanding of efficiency gains beyond existing statistical paradigms.
5 Simulation studies
This section demonstrates the performance of the proposed procedures through simulations. We first introduce the settings of parameter generation and then present results.
In Section 4, we highlight that the proposed estimators do not require the columns of $Z^{\star}$ and $W_{t}^{\star}$ to be nearly orthogonal, characterized through $\mathcal{G}([Z^{\star},W^{\star}])$. To demonstrate that, we next generate true parameters satisfying $\mathcal{G}([Z^{\star},W^{\star}])=\Omega\otimes\mathrm{I}_{k}/(2\sqrt{k})$, where $\Omega=(\omega_{i,l})\in\mathbb{R}^{(1+T)\times(1+T)}$ with $\omega_{i,i}=1$, and $1/(2\sqrt{k})$ is a normalization scalar. This Gram structure implies that columns of $Z^{\star}$ are orthogonal, columns of $W_{t}^{\star}$ are orthogonal, and $\cos(Z_{j}^{\star},W_{t,j}^{\star})=\omega_{1,1+t}$ and $\cos(W_{t,j}^{\star},W_{s,j}^{\star})=\omega_{1+t,1+s}$, where $\cos(a,b)=\langle a,b\rangle/(\|a\|_{2}\|b\|_{2})$. When $\omega_{1,1+t}$ is larger, the two vectors $Z_{j}^{\star}$ and $W_{t,j}^{\star}$ are less orthogonal; similar interpretation applies to $\omega_{1+t,1+s}$. We next consider three cases (A)–(C) with $\Omega$ set as follows:
(A) (B) (C)
$\Omega$ $\begin{pmatrix}1&\mathbf{0}\\ \mathbf{0}&\mathrm{I}_{T}\end{pmatrix}$ $\begin{pmatrix}1&\phi\mathbf{1}_{T}^{\top}\\ \phi\mathbf{1}_{T}&\Sigma_{T}(\rho)\end{pmatrix}$ $\begin{pmatrix}\mathrm{I}_{1+T_{o}}&\mathbf{0}\\ \mathbf{0}&\Sigma_{T-T_{o}}(1)\end{pmatrix}$,
[…]
In each setting, we evaluate the performance of the proposed Algorithms 1 and 2, referred to as SS-Hunting and SS-Refinement, respectively. The two methods MultiNeSS and MultiNeSS+ in MacDonald et al., (2022) are compared as our models are similar. Their errors are not presented under Poisson distribution as codes are unavailable. We next focus on illustrating the error rate of estimating $Z^{\star}$ with respect to $T$, highlighting the improvement over using a single network. The error rates for estimating $W_{t}^{\star}$’s are not improved over $T$, which is consistent with our expected oracle rates in Section 2.3, and detailed results are deferred to Section G of the Supplementary Material. For the ease of visualization, we vary $T\in\{5,10,20,40,80\}$ while fixing $n=400$ and $k_{t}=k=2$ for $t=1,\ldots,T$. For a fair comparison with MacDonald et al., (2022), we present error $\|\widehat{ZZ^{\top}}-Z^{\star}Z^{\star\top}\|_{\mathrm{F}}^{2}/n$, equivalent to $\operatorname{dist}^{2}(\hat{Z},Z^{\star})$ up to constants multiplied (Tu et al.,, 2016).
(a) Bernoulli distribution
(b) Gaussian distribution
(c) Poisson distribution
Figure 1: Empirical estimation errors versus $T$ under Case (A). Lines connect median estimation errors over 100 repetitions, and error bars are obtained from the 0.05 and 0.95 quantiles of those repetitions. Axes are in the log scale.
Figures 13 present empirical estimation errors of the four estimators under Cases (A)–(C), respectively. In Figure 1, all the errors are inverse proportional to $T$, achieving the oracle error rate discussed in Section 2.3, and SS-Refinement and MultiNeSS+ achieve the smallest errors. In Figure 2, SS-Refinement achieves the smallest error across three distributions, while SS-Hunting performs similarly under Gaussian and Poisson distributions but not Bernoulli distribution. Different from Figure 1, errors of MultiNeSS and MultiNeSS+ are no longer inverse proportional to $T$. Under Case (C), errors of MultiNeSS and MultiNeSS+ explode and are not presented in Figure 3 for visual clarity. More details of their results are given in Section G.3 of the Supplementary Material. Unlike in Figures 1 and 2, errors of SS-Hunting are not inverse proportional to $T$, whereas SS-Refinement still achieves that, demonstrating efficiency gain from the likelihood refinement.
(a) Bernoulli distribution
(b) Gaussian distribution
(c) Poisson distribution
Figure 2: Empirical estimation errors versus $T$ under Case (B). (Similar to Figure 1.)
(a) Bernoulli distribution
(b) Gaussian distribution
(c) Poisson distribution
Figure 3: Empirical estimation errors versus $T$ under Case (C). (Similar to Figure 1.)
Intuitively, Cases (A)–(C) correspond to more and more challenging scenarios where the relationship between latent spaces become non-orthogonal and more complex. The non-ideal performances of MultiNeSS and MultiNeSS+ under Cases (B) and (C) align with Remark 4 showing that our methods would require less restriction on the latent spaces. For our two proposed estimators, SS-Refinement achieves the desired inverse proportional to $T$ error rate across all cases and distribution, whereas SS-Hunting fails to achieve that in challenging Case (C). This is consistent with our theoretical results in Theorems 1 and 2 providing general error rates under weak assumptions about latent spaces. The results demonstrate the effectiveness of the proposed efficient estimation strategy.
6 Data analysis
We analyze a multiple-network dataset of lawyers (Lazega,, 2001), which contains three types of connection relationships: coworker, friendship, and advice, between seventy-one lawyers in a US corporate law firm. Within each type of network, we consider binary and undirected edges between lawyers indicating whether connections of each type exist between them. Since there is no ground truth of latent spaces in practice, we compare the estimated latent embeddings with observed node-wise features for interpretation.
As network edges are binary, we adopt Bernoulli distribution with logistic link and estimate latent dimensions by Algorithm B.1, which gives $\hat{k}_{1}=6$, $\hat{k}_{2}=4$, $\hat{k}_{3}=1$, and $\hat{k}=2$. Then we estimate latent embedding vectors under estimated latent dimensions by SS-Refinement. As a comparison, we also apply MultiNeSS+ in MacDonald et al., (2022) and MASE in Arroyo et al., (2021) with the dimension of the shared space set to be two.
(a) $S$-Refinement
(b) MultiNeSS+
(c) MASE
Figure 4: Estimated shared latent vectors $z_{i}$’s using three methods.
We first illustrate the shared latent embeddings estimated by the three methods. Figure 4 (a) and (b) present the leading two principal component scores of the shared latent component $Z$ estimated by SS-Refinement and MultiNeSS+, respectively. Figure 4 (c) displays latent embeddings estimated by MASE that are rotated to approximate (a) and (b) for the ease of comparison. In Figure 4, each point is colored according to the office location of the corresponding lawyer. The results of all three methods show that the shared latent embeddings are correlated with the office locations of lawyers. This suggests that office locations could play a shared role in forming multiple types of connections between lawyers.
(a) $S$-Refinement
(b) MultiNeSS+
Figure 5: Estimated individual latent vectors $w_{1,i}$’s using two methods.
We next compare the individual latent embeddings $W_{t}$’s estimated by SS-Refinement and MultiNeSS+, where we point out that MASE does not directly give comparable estimates and thus is not included. Figures 5 and 6 present the estimated individual latent component after projected to its leading two principal components for the coworker and friendship networks, respectively; see more details in Section H.2 of the Supplementary Material. In Figure 5, points are colored according to lawyers’ practice. The results show that latent vectors estimated by both two methods demonstrate association with lawyars’ practice, whereas the separation between litigation and corporate lawyers appears to be more obvious in estimates given by SS-Refinement. This discrepancy might be attributed to the fact that MultiNeSS+ tends to encourage the column spaces of $Z$ and $W_{t}$’s to be orthogonal, which is not required in our approach and thus leads to distinct estimates in practice. In Figure 6, points are colored based on the lawyers’ status. Similarly to Figure 5, latent vectors estimated by both two methods exhibit correlation with a common covariate, lawyers’ status, even though the patterns of points are not exactly the same.
(a) $S$-Refinement
(b) MultiNeSS+
Figure 6: Estimated individual latent vectors $w_{2,i}$’s using two methods.
In summary, our analysis shows that investigating shared and individual latent embeddings reveals key underlying factors driving the formation of heterogeneous multiplex networks. The proposed analysis paradigm could offer us a deeper understanding of multifaceted relationships and unveil inherent structures of interlinked systems across various domains.
7 Discussions
This work establishes a general analysis framework for efficient estimation of node-wise latent embeddings in multiple heterogeneous networks. It shows that aggregating multiplex networks can improve the efficiency of estimating shared latent spaces. In particular, we develop a novel two-stage approach: the first stage hunts for shared latent embedding spaces, and the second stage improves the statistical efficiency with likelihood information. We establish estimation error bounds for the estimators of the two stages. The final estimator has been shown to achieve the oracle rates of component for both the shared and individual components according to their dimensions. The new developments would provide statistical insights into how to efficiently aggregate heterogeneous datasets and unravel fundamental structures in complex interconnected systems.
The new developments pave the way for a broad range of future studies. First, accurate estimation of latent embedding spaces plays a crucial role in subsequent analytical tasks including, but not limited to, understanding latent confounders across multifaceted interconnected systems, adjusting dependence of observations for causal inference (McFowland III and Shalizi,, 2023; Nath et al.,, 2024; Hayes et al.,, 2024), and prediction with or on noisy network data (Ma et al.,, 2020; Le and Li,, 2022; Lunde et al.,, 2023).
Second, another interesting future research direction is to extend the analysis to various other network models. For instance, to analyze networks exhibiting significant degree heterogeneity, we could add network-specific and node-specific degree heterogeneity parameters in current model, similarly to Zhang et al., (2020) and He et al., (2025). Moreover, other forms of interactions could be considered, such as using a general kernel function beyond inner product (Rubin-Delanchy et al.,, 2022; MacDonald et al.,, 2022), or modeling directed network edges with asymmetric latent embeddings (Perry and Wolfe,, 2013; Yan et al.,, 2019). In addition, when there are observed covariates (Yan and Sarkar,, 2021; Huang et al.,, 2024), it would be worthwhile to devise appropriate procedures for covariate adjustments. The methods and theoretical tools developed in this work provide a versatile toolbox that could be useful under diverse scenarios of network analysis.
Finally, our current analysis of efficient estimation requires the likelihood to be correctly specified. However, in many applications, distributions of noisy networks may deviate from their prespecified forms. This misspecification might alter the interpretation of latent embeddings and impede efficient estimations. Understanding the effect of distribution misspecification would be an important future research direction. Such studies could provide us better understanding towards underlying data structures and lead to reliable and robust conclusions (Rubin-Delanchy,, 2020).
Supplementary Material
Due to space limitation, additional details, including individual estimation in Section 3.1, estimating latent dimensions, and proofs are deferred to the Supplementary Material.
Acknowledgment
We are grateful to the editors and two anonymous reviewers for their valuable comments. YH was partially supported by NSF Grant DMS-2515523 and Wisconsin Alumni Research Foundation.